OpenStudy (anonymous):

Find a cubic function with the given zeros. Square root of two., - Square root of two., -2 f(x) = x3 + 2x2 - 2x + 4 f(x) = x3 + 2x2 + 2x - 4 f(x) = x3 - 2x2 - 2x - 4 f(x) = x3 + 2x2 - 2x - 4 @amistre64

OpenStudy (anonymous):

Please help! @robtobey

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x) = 5(x + 7)2(x - 7)3 -7, multiplicity 2; 7, multiplicity 3 4, multiplicity 1; -7, multiplicity 3; 7, multiplicity 3 -7, multiplicity 3; 7, multiplicity 2 4, multiplicity 1; 7, multiplicity 1; -7, multiplicity 1

OpenStudy (anonymous):

@iGreen pleaase help!!

OpenStudy (mathstudent55):

For the first question above, a polynomial function with roots x1, x2, x3, ... has equation (x - x1)(x - x2)(x - x3)... = 0 Write all the binomial factors using the roots you were given, then multiply them out.

OpenStudy (mathstudent55):

For the second question, from the choices it is obvious some of those numbers are exponents. Is the equation \(f(x) = 5(x + 7)^2(x - 7)^3\) If so, a factor of x + 7 will give a root of x = -7 since x + 7 = 0 x = -7 A factor of x - 7 give x = 7 since x - 7 = 0 x = 7 The multiplicity comes from the exponent in each factor. Since the factor x + 7 has an exponent of 2, x = -7 has multiplicity 2. Since the factor x - 7 has exponent 3, x = 7 has multiplicity 3.

OpenStudy (mathstudent55):

Think of it this way: \(f(x) = 5(x + 7)(x + 7)(x - 7)(x - 7)(x - 7)\) Set the function equal to zero: \(5(x + 7)(x + 7)(x - 7)(x - 7)(x - 7) = 0\) Now set each factor that has x in it equal to zero: \(x + 7 = 0\) or \(x + 7 = 0\) or \(x - 7 = 0\) or \(x - 7 = 0\) or \(x - 7 = 0\) Solve all the equations: x = -7 or x = -7 or x = 7 or x = 7 or x = 7 or x = 7 Now count how many roots are -7 and how many are 7: 2 roots -7 3 roots 7 Answer is: -7, multiplicity 2; 7, multiplicity 3

OpenStudy (anonymous):

Thank you so much! @mathstudent55

OpenStudy (mathstudent55):

You're welcome.